As the price of tickets rises from $200 to

$250, what is the price elasticity of demand

for (i) business travelers and (ii) vacationers?

(Use the midpoint method in your

calculations.)

b. Why might vacationers have a different

elasticity from business travelers?


1
Expert's answer
2021-10-04T14:10:15-0400

Elasticity of demand depicts how much a consumer responds to change in price level.

(a)

"Ed=\frac{\frac{Q_2-Q_1}{\frac{Q_2+Q_1}{2}}}{\frac{P_2-P_1}{\frac{P_2+P_1}{2}}}"


(i)

For business travelers


"Ped,Ed=\frac{ \frac{2000-2100}{\frac{2000+2100}{2}}}{\frac{200-150}{\frac{200+150}{2}}}"


"Ed=-0.17"

"|Ed|=0.17<1; inelastic"



(ii)

"Ped,Ed=\frac{ \frac{800-1000}{\frac{800+1000}{2}}}{\frac{200-150}{\frac{200+150}{2}}}"



"Ed=-0.78"

"|Ed|=0.78<1;inelastic"

b) Vacationers are casual with their plan whereas business travelers are traveling with business purpose thus their elasticity of demand for tickets is less affected with change in price.